OpenStudy (anonymous):

how do you find the center for a circle with the equation x^2+y^2=5x+6y

OpenStudy (raden):

general form equation of circle is : x^2+y^2+Ax+By+c=0 with the center is C(a,b) : a=-1/2*A, b=-1/2*B

OpenStudy (anonymous):

is the center -5,6? can you explain the last part?

OpenStudy (raden):

ok, given equation of circle : x^2+y^2=5x+6y or x^2+y^2-5x-6y=0 so, A=-5 and B=-6 to get its center, use formula like i said before a = -1/2*A = -1/2 * (-5) = ....? b = -1/2*B = -1/2 * (-6) = ....?

OpenStudy (anonymous):

alright so is it a=2.5 and b= 3 ?

OpenStudy (raden):

thats right

OpenStudy (anonymous):

so the radius is 0? how would you draw this?

OpenStudy (raden):

to get its radius, u can use formula :